# Quantum Doubly Stochastic Transformers

### v1

@@ 4 Theoretical result on number of DSMs
The optimal way of studying the expressivity of a doubly-stochastic operator empirically would enumerate all DSMs in a given Birkhoff polytope $\Omega_{n}$ and assess for each DSM whether it can be reached (or how closely). The exact volume of the Birkhoff polytope is an open problem in mathematics [10] which limits our ability to study expressivity theoretically. In practice, one can assume a certain discretization $p\in\mathbb{N}_{+}\text{ s.t., }\bm{\mathrm{P}}_{ij}\in\{0,\frac{1}{p-1},...,1\}$, e.g., if $p=2$ then $\bm{\mathrm{P}}_{ij}\in\{0,1\}$. In that case there are $n!$ DSMs. In Appendix D we provide a partial derivation for the combinatorial problem of identifying the function $f(n,p)\rightarrow\mathbb{N}$ returning the number of DSMs. The basic idea is as follows. Given that a $n\times n$ DSM has $(n-1)^{2}$ degrees of freedom, there are $p^{(n-1)^{2}}$ candidate matrices. Not all of these can be turned into DSMs because of two constaints, (1) the sum of any row or column must not exceed 1 and (2) the sum of the $n-1\times n-1$ submatrix must not be below $n-2$ [10]. This allows to decompose $f$ into $f(n,p)=p^{(n-1)^{2}}-c_{1}-c_{2}+c_{12}$ where $c_{1}$ and $c_{2}$ measure how often the constraints are violated and $c_{12}$ corrects for the cases where both constraints are violated. For details see Appendix D.
@@ 5.1 Experimental Setup
We evaluate different flavors of DSFormers obtained through replacing the Softmax function with any of the DSM operators described above. When integrating our QontOT-derived operator into a ViT we obtain our hybrid quantum-classical doubly stochastic Transformer (QDSFormer). In the following, we refer to QontOT as the attention flavor which contains the quantum circuit whereas QDSFormer denotes, more broadly, any Transformer with quantum doubly-stochastic attention. To date, the only realization of a QDSFormer is through QontOT. Among all operators, the classical ones are non-parametric whereas this quantum operator can be optimized during training. Therefore, one could theoretically optimize circuit parameters concurrently with Transformer training. However, the ViTs we study contains up to $4$ attention layers, with a batch size of $512$, yielding $2048$ samples to optimize in a single forward pass. We predict, unless mentioned otherwise, $8\times 8$ DSMs, use $16$ circuit layers and $4$ auxiliary qubits ($16$ qubits in total). Running the circuit on quantum hardware requires $\Omega(n^{2}/\varepsilon^{2})$ shots to obtain satisfactory sampling error [37]. Assuming a precision of $\varepsilon=0.01$, this is in the order of $640$k shots per sample. Since quantum hardware operates on kHz frequency, execution and online optimization on hardware is unfortunately not (yet) feasible. Therefore, we perform exact statevector simulation with Qiskit [26] and implement three circuit training strategies:
@@
    Differentiable: Joint optimization of circuit and Transformer parameters through backpropagation, akin to integrating the circuit as a neural network layer. This is by far the slowest given the difficulty of gradient propagation through quantum circuits [2].
@@
    Mixed: A mixed strategy where Transformer training is interleaved with $200$ steps of gradient-free circuit optimization with Nevergrad [6] on a per-epoch basis.
@@
    Static: The circuit is used in pure inference mode with parameters obtained from a 24-qubit DSM prediction experiment on quantum hardware [37].
From the operators studied in Section 3.1, we discard the Birkhoff projection due to its non-differentiability and low DSM-diversity (Figure 2). For comparison, we further include the NormSoftmax [27], here denoted as Softmaxσ, that attenuates attention by taking the minimum of the empirical and the expected standard deviation $\tau:=\sqrt{d_{k}}$:
$\bm{\mathrm{A}}=\text{Softmax}\left(\frac{\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}}{\min{\left(\sigma(\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}),\tau\right)}}\right).$ (6)
@@ 5.2 Data sets
We evaluate the different ViTs on MNIST [33], Fashion MNIST [57], seven datasts from the MedMNIST benchmark [58] and a compositional task requiring multistep reasoning as proposed by Hoffmann et al. [24]. In that task, a $2\times 2$ grid contains two MNIST digits (upper left and lower right) and two FashionMNIST items (upper right and lower left). If the digits have equal value, the label is the upper right fashion item, otherwise it is the bottom left fashion item. Performance typically ramps up quickly to $\sim 50\%$ because the model learns to attend one (and only one) of the FashionMNIST images. Upon continued training with a long saturation phase, a ViT suddenly grasps the relationship of the MNIST digits to the classification task and then climbs rapidly to a $90-95\%$ accuracy. The moment of abrupt improvement is called ”Eureka moment” [24]. The dataset is split into $60$K ($10$K) training (validation) examples. To accommodate the $8\times 8$ attention matrix, each image from MNIST, FashionMNIST and MedMNIST is split into $7$ horizontal stripes and a CLS token is pre-pended to the patch sequence. For the Eureka dataset we use a patch size of $14\times 28$ pixels and mean-pooling.
@@ 5.3 Empirical results
This confirms the finding that doubly stochastic attention can improve ViTs [42]. Moreover, in both cases, adding more circuit layers increases performance logarithmically and with $4$ or $8$ circuit layers the ViT performance is surpassed. Exact numerical results are provided in Appendix Table A3.
@@
Table 1: $L$-layered ViT validation accuracy on FashionMNIST for different attention methods. QontOT uses $16$ circuit layers. Mean/std computed from $5$ trainings.
@@
1 $\underline{86.5}_{\pm 0.19}$ $75.3_{\pm 4.61}$ $\textbf{87.1}_{\pm 0.26}$ $85.6_{\pm 0.10}$ $84.2_{\pm 3.64}$
2 $88.9_{\pm 0.12}$ $84.6_{\pm 2.11}$ $\underline{89.3}_{\pm 0.07}$ $\textbf{90.0}_{\pm 0.15}$ $89.1_{\pm 0.73}$
3 $\underline{89.4}_{\pm 0.28}$ $86.3_{\pm 2.69}$ $\underline{89.4}_{\pm 0.11}$ $\textbf{90.3}_{\pm 0.13}$ $\underline{89.4}_{\pm 0.77}$
4 $\underline{89.7}_{\pm 0.29}$ $87.1_{\pm 1.15}$ $89.5_{\pm 0.07}$ $\textbf{90.3}_{\pm 0.14}$ $89.1_{\pm 1.08}$
The same result was obtained on MNIST and, this time, QontOT outperformed softmaxed attention also for one ViT layer (see Table 2).
Table 2: Identical to Table 1 but for MNIST.
[…]
1 $89.1_{\pm 12.5}$ $66.7_{\pm 22.5}$ $\textbf{96.6}_{\pm 0.10}$ $93.9_{\pm 0.11}$ $\underline{94.3}_{\pm 1.97}$
2 $98.1_{\pm 0.33}$ $93.0_{\pm 4.57}$ $\underline{98.3}_{\pm 0.13}$ $\textbf{98.4}_{\pm 0.05}$ $98.2_{\pm 0.27}$
3 $\underline{98.6}_{\pm 0.11}$ $97.7_{\pm 0.65}$ $\underline{98.6}_{\pm 0.13}$ $\textbf{98.7}_{\pm 0.06}$ $\underline{98.6}_{\pm 0.12}$
4 $\textbf{98.8}_{\pm 0.10}$ $97.9_{\pm 0.71}$ $\underline{98.7}_{\pm 0.11}$ $\textbf{98.8}_{\pm 0.07}$ $97.9_{\pm 1.57}$
In further experiments with more ViT layers performance assimilated and plateaued due to the simplicity of the datasets. But we saw scant further improvement for more than $16$ circuit layers. For a barplot visualization of Table 1 and 2 see Appendix Figure A4/A5. Notably, QontOT offers great flexibility in the type of ansatz for the quantum circuit [37]. We observed only minor differences between four different ansatz types, with three of them outperforming the ViT, underlining the generality of the finding (Appendix Table A1).
A compelling finding is that the static version of the QontOT-attention did perform as good or even better than the optimized one (see Appendix Figure A7). We tested an end-to-end optimizable QDSFormer where circuit and ViT parameters are jointly optimized. Such end-to-end training is not only slower, but also had lower accuracy than the static configuration, for both MNIST and FashionMNIST and $1$ and $8$ circuit layers (Figure A7). This may be caused by Barren plateaus [38] (i.e., gradients are largely constant along most directions), a widespread phenomenon in variational quantum circuits that slows down learning. We further experimented with a ”mixed” training strategy where the circuit is trained every $n$-th epoch. This did not reveal a clear benefit for more frequent circuit optimization (see Appendix Figure A6), potentially due to higher volatility of the circuit. We therefore use the static, faster configuration in all remaining experiments.
Table 3: Test accuracy for different MedMNIST datasets across five attention types in a 2-layer ViT. QontOT uses 16 circuit layers.
@@
Next, we repeated the comparison to the four classical attention types on larger datasets (up to $240$k images) from the MedMNIST benchmark [58]. In $5$ out of $7$ datasets, the QDSFormer obtained significantly better results than all other methods (Table 3), with a mean accuracy increase of $1.3\%$ compared to a standard ViT. Notably, none of the other attention types can improve upon the standard ViT.
Another important advantage of the QDSFormer is its stabilizing effect. Among repeated training runs, the performance variation (i.e., test accuracy variance) is consistently lower than for all classical methods (e.g., Table 2 or Table 3). Notably, with 1 ViT-layer and softmax-based attention some trainings on MNIST failed to converge.
@@
Figure 5: Validation accuracy for each training epoch, highlighting the abrupt learning referred to as Eureka Moment (EM) on the compositional dataset. Confidence bounds from 5 runs.
This major improvement, achieved with an extremely lightweight quantum circuit (1 layer) also consolidated across different learning rates (Table 4).
Table 4: Results on the Eureka dataset across different models and learning rates. EM@Ep denotes the average epoch of the EM, for runs without EM, the epoch is considered to be $100$.
@@
$1$e-$3$ Acc. $53.4_{\pm 0.09}$ $48.6_{\pm 1.12}$ $49.0_{\pm 0.85}$ $\underline{61.2}_{\pm 11.1}$ $\textbf{70.0}_{\pm 13.9}$
EM@Ep $\textbf{72.2}_{\pm 14.2}$ $\underline{74.8}_{\pm 8.43}$
@@
$7$e-$4$ Acc. $53.9_{\pm 0.16}$ $50.9_{\pm 0.83}$ $50.9_{\pm 0.65}$ $\underline{72.9}_{\pm 10.6}$ $\textbf{82.3}_{\pm 14.4}$
EM@Ep $\underline{57.7}_{\pm 10.6}$ $\textbf{43.3}_{\pm 7.01}$
@@
$5$e-$4$ Acc. $61.1_{\pm 14.6}$ $51.0_{\pm 0.52}$ $51.6_{\pm 0.24}$ $\underline{66.4}_{\pm 15.6}$ $\textbf{89.4}_{\pm 0.09}$
EM@Ep $61.0$ $\underline{72.6}_{\pm 21.6}$ $\textbf{43.6}_{\pm 8.25}$
# EM $1/5$ $0/5$ $0/5$ $\underline{\vphantom{y}\smash{2/5}}$ 5/5
Next, to assess the potential of the QDSFormer beyond image data and simultaneously study the scalability of the QDSFormer, we applied it on a novel dataset for time-series classification of InfraRed (IR) spectra of molecules into $37$ functional groups [4]. This dataset contains almost $1$M samples and we scale up the circuit to produce $4$x larger attention matrices ($16\times 16$). On this dataset, the performance differences are marginal and the QDSFormer performs on par with a standard ViT (for details se Appendix Table A2). This shows that the QDSFormer can meaningfully generalize to domains like scientific data and tasks like multi-label classification.
Since all above results were obtained through statevector simulation, we conducted a final experiment to understand the detrimental effect of quantum noise induced by real quantum hardware (via the publicly available IBM Quantum Platform, for details see Appendix E). This 14-qubit experiment tests the potential for a hybrid hardware training. Despite various light error mitigation techniques, the obtained doubly stochastic attention matrices consistently show high entropy (i.e., a tendency toward more uniform distributions), even for larger shot counts.
Figure 6: Hardware experiment on different quantum computers available via IBM Quantum Platform. A The Frobenius distance between the hardware-obtained DSM to its noise-free equivalent. B: The spearman rank correlation between the $64$ values in the noise-free and hardware-obtained DSMs show that the ordering of values is largely preserved. D: Execution time in seconds on different quantum computers.
However, as Figure 6B shows, when comparing to the noise-free ground truth attention matrix, the ordering of the values in the attention matrix was preserved with high precision (spearman $\rho>0.9$ even for moderate shot count). Since the circuit runs within a ViT, successfully preserving the ordering will be key (to not destroy signal). Instead, numerical exactness (cf. Figure 6A) may be compromised and we suspect that embedding such a noisy quantum attention block (which preserves peak attention scores but also increases entropy) into a Transformer could even be advantageous. The additional entropy may avoid vanishing gradients and act as a form of regularization. This effect is particularly notable compared to its noise-free analog, which remains classically intractable if sufficient qubits are used.

### v2

@@ 4 Theoretical result on number of DSMs
The optimal way of studying the expressivity of a doubly-stochastic operator empirically would enumerate all DSMs in a given Birkhoff polytope $\Omega_{n}$ and assess for each DSM whether it can be reached (or how closely). The exact volume of the Birkhoff polytope is an open problem in mathematics [chan1999volume] which limits our ability to study expressivity theoretically. In practice, one can assume a certain discretization $p\in\mathbb{N}_{+}\text{ s.t., }\bm{\mathrm{P}}_{ij}\in\{0,\frac{1}{p-1},...,1\}$, e.g., if $p=2$ then $\bm{\mathrm{P}}_{ij}\in\{0,1\}$. In that case there are $n!$ DSMs. In Appendix E we provide a partial derivation for the combinatorial problem of identifying the function $f(n,p)\rightarrow\mathbb{N}$ returning the number of DSMs. The basic idea is that a $n\times n$ DSM has $(n-1)^{2}$ degrees of freedom, thus there are $p^{(n-1)^{2}}$ candidate matrices. Not all of these can be turned into DSMs because of two constaints, (1) the sum of any row or column must not exceed 1 and (2) the sum of the $n-1\times n-1$ submatrix must not be below n2n-2 [chan1999volume]. This allows to decompose $f$ into $f(n,p)=p^{(n-1)^{2}}-c_{1}-c_{2}+c_{12}$ where $c_{1}$ and $c_{2}$ measure the constraint violations and $c_{12}$ discounts cases where both constraints are violated. For details see Appendix E.
@@ 5.1 Experimental Setup
We evaluate different flavors of DSFormers obtained through replacing the Softmax function with any of the DSM operators described above. When integrating our QontOT-derived operator into a ViT we obtain our hybrid quantum-classical doubly stochastic Transformer (QDSFormer). In the following, we refer to QontOT as the attention flavor which contains the quantum circuit whereas QDSFormer denotes, more broadly, any Transformer with quantum doubly-stochastic attention. To date, the only realization of a QDSFormer is through QontOT. Among all operators, the classical ones are non-parametric whereas this quantum operator can be optimized during training. Therefore, one could theoretically optimize circuit parameters concurrently with Transformer training. However, the ViTs we study contains up to $4$ attention layers, with a batch size of $512$, yielding $2048$ samples to optimize in a single forward pass. We predict, unless mentioned otherwise, $8\times 8$ DSMs, use $16$ circuit layers and $4$ auxiliary qubits ($16$ qubits in total). Running the circuit on quantum hardware requires $\Omega(n^{2}/\varepsilon^{2})$ shots to obtain satisfactory sampling error [mariella2024quantum]. Assuming a precision of $\varepsilon=0.01$, this is in the order of 640640k shots per sample. Since quantum hardware operates on kHz frequency, execution and online optimization on hardware is unfortunately not (yet) feasible. Therefore, we perform exact statevector simulation with Qiskit [javadi2024quantum] and implement three circuit training strategies:
@@
    Differentiable: Joint optimization of circuit and Transformer parameters through backpropagation, akin to integrating the circuit as a neural network layer. This is by far the slowest given the difficulty of gradient propagation through quantum circuits [abbas2024quantum].
@@
    Mixed: A mixed strategy where Transformer training is interleaved with $200$ steps of gradient-free circuit optimization with Nevergrad [bennet2021nevergrad] on a per-epoch basis.
@@
    Static: The circuit is used in pure inference mode with parameters obtained from a 24-qubit DSM prediction experiment on quantum hardware [mariella2024quantum].
@@ 5.2 Data sets
We evaluate all ViTs on MNIST [lecun2010mnist], Fashion MNIST [xiao2017fashionmnistnovelimagedataset], seven datasets from the MedMNIST benchmark [medmnistv1] and a compositional task requiring multistep reasoning [hoffmann2024eureka]. In that task, a $2\times 2$ grid contains two MNIST digits (upper left and lower right) and two FashionMNIST items (upper right and lower left). If the digits have equal value, the label is the upper right fashion item, otherwise it is the bottom left fashion item. Performance typically ramps up quickly to $\sim 50\%$ because the model learns to attend one (and only one) of the FashionMNIST images. Upon continued training with a long saturation phase, a ViT suddenly grasps the relationship of the MNIST digits to the classification task and then climbs rapidly to a $90-95\%$ accuracy. The moment of abrupt improvement is called "Eureka moment" [hoffmann2024eureka]. The dataset is split into 6060K (1010K) training (validation) examples. To accommodate the $8\times 8$ attention matrix, each image from MNIST, FashionMNIST and MedMNIST is split into $7$ horizontal stripes and a CLS token is pre-pended to the patch sequence. For the Eureka dataset we use a patch size of $14\times 28$ pixels and mean-pooling.
@@ 5.3 Empirical results
Table 1: Validation accuracy of $L$-layered ViT on FashionMNIST and MNIST for different attention methods. QontOT uses $16$ circuit layers. Mean/std computed from 55 trainings.
FashionMNIST MNIST
@@
1 $\underline{86.5}_{\pm 0.2}$ $75.3_{\pm 4.6}$ $\textbf{87.1}_{\pm 0.3}$ $85.6_{\pm 0.1}$ $84.2_{\pm 3.6}$ $89.1_{\pm 12.5}$ $66.7_{\pm 22.5}$ $\textbf{96.6}_{\pm 0.1}$ $93.9_{\pm 0.1}$ $\underline{94.3}_{\pm 2.0}$
2 $88.9_{\pm 0.1}$ $84.6_{\pm 2.1}$ $\underline{89.3}_{\pm 0.1}$ $\textbf{90.0}_{\pm 0.2}$ $89.1_{\pm 0.7}$ $98.1_{\pm 0.3}$ $93.0_{\pm 4.6}$ $\underline{98.3}_{\pm 0.1}$ $\textbf{98.4}_{\pm 0.1}$ $98.2_{\pm 0.3}$
3 $\underline{89.4}_{\pm 0.3}$ $86.3_{\pm 2.7}$ $\underline{89.4}_{\pm 0.1}$ $\textbf{90.3}_{\pm 0.1}$ $\underline{89.4}_{\pm 0.8}$ $\underline{98.6}_{\pm 0.1}$ $97.7_{\pm 0.7}$ $\underline{98.6}_{\pm 0.1}$ $\textbf{98.7}_{\pm 0.1}$ $\underline{98.6}_{\pm 0.1}$
4 $\underline{89.7}_{\pm 0.3}$ $87.1_{\pm 1.2}$ $89.5_{\pm 0.1}$ $\textbf{90.3}_{\pm 0.1}$ $89.1_{\pm 1.1}$ $\textbf{98.8}_{\pm 0.1}$ $97.9_{\pm 0.7}$ $\underline{98.7}_{\pm 0.1}$ $\textbf{98.8}_{\pm 0.1}$ $97.9_{\pm 1.6}$
In further experiments with more ViT layers performance assimilated and plateaued due to the simplicity of the datasets. But we saw scant further improvement for more than $16$ circuit layers. For a barplot visualization of Table 1 see Appendix Figure A5/A6. Notably, QontOT offers great flexibility in the type of ansatz for the quantum circuit [mariella2024quantum]. We observed only minor differences between four different ansatz types, with three of them outperforming the ViT, underlining the generality of the finding (Appendix Table A2). A compelling aspect is that the static version of the QontOT-attention did perform as good or even better than the optimized one (see Appendix Figure A8). We tested an end-to-end optimizable QDSFormer where circuit and ViT parameters are jointly optimized. Such end-to-end training is not only slower, but also had lower accuracy than the static configuration, for both MNIST and FashionMNIST and $1$ and $8$ circuit layers (Figure A8). This may be caused by Barren plateaus [mcclean2018barren] (i.e., gradients are largely constant along most directions), a widespread phenomenon in variational quantum circuits that slows down learning. We further experimented with a "mixed" training strategy where the circuit is trained every $n$-th epoch. This did not reveal a clear benefit for more frequent circuit optimization (see Appendix Figure A7), potentially due to higher volatility of the circuit. We therefore use the static, faster configuration in all remaining experiments.
Table 2: Test accuracy for MedMNIST datasets across 55 attention types in a 2-layer ViT.
@@
Next, we repeated the comparison to the four classical attention types on larger datasets (up to 240240k images) from the MedMNIST benchmark [medmnistv1]. In $5$ out of $7$ datasets, the QDSFormer obtained significantly better results than all other methods (Table 2), with a mean accuracy increase of $1.3\%$ compared to a standard ViT. Notably, none of the other attention types can improve upon the standard ViT. Another important advantage of the QDSFormer is its stabilizing effect. Among repeated training runs, the performance variation (i.e., test accuracy variance) is consistently lower than for all classical methods (e.g., Table 1 or Table 2). Notably, with 1 ViT-layer and softmax-based attention some trainings on MNIST failed to converge.
@@
Next, to assess the potential of the QDSFormer beyond image data and simultaneously study the scalability of the QDSFormer, we applied it on a novel dataset for time-series classification of InfraRed (IR) spectra of molecules into $37$ functional groups [alberts2025unraveling]. This dataset contains almost 11M samples and we scale up the circuit to produce 44x larger attention matrices ($16\times 16$). On this dataset, the performance differences are marginal and the QDSFormer performs on par with a standard ViT (for details see Appendix Table A3). This shows that the QDSFormer can meaningfully generalize to domains like scientific data and tasks like multi-label classification.
@@
11e-33 Acc. $53.4_{\pm 0.1}$ $48.6_{\pm 1.1}$ $49.0_{\pm 0.9}$ $\underline{61.2}_{\pm 11.0}$ $\textbf{70.0}_{\pm 13.9}$
EM@Ep $\textbf{72.2}_{\pm 14.2}$ $\underline{74.8}_{\pm 8.4}$
@@
77e-44 Acc. $53.9_{\pm 0.2}$ $50.9_{\pm 0.8}$ $50.9_{\pm 0.7}$ $\underline{72.9}_{\pm 10.0}$ $\textbf{82.3}_{\pm 14.0}$
EM@Ep $\underline{57.7}_{\pm 11.0}$ $\textbf{43.3}_{\pm 7.0}$
@@
55e-44 Acc. $61.1_{\pm 15.0}$ $51.0_{\pm 0.5}$ $51.6_{\pm 0.2}$ $\underline{66.4}_{\pm 16.0}$ $\textbf{89.4}_{\pm 0.1}$
EM@Ep $61.0$ $\underline{72.6}_{\pm 22.0}$ $\textbf{43.6}_{\pm 8.3}$
# EM $1/5$ $0/5$ $0/5$ $\underline{2/5}$ 5/5
(a) Results on the Eureka dataset across different models and learning rates. EM@Ep denotes the average epoch of the EM; runs without EM are set to epoch 100.
(b) Validation accuracy per epoch, highlighting the Eureka Moment on the compositional dataset. Confidence bounds from 5 runs.
Figure 4: (a) Eureka results across attention methods. (b) QDSFormer antedates the Eureka Moment (EM).
Since all above results were obtained through statevector simulation, we conducted a final experiment to understand the detrimental effect of quantum noise induced by real quantum hardware via the publicly available IBM Quantum Platform. We used the three machines Torino (Heron R1, 133 qubits, error per layered gate: $1.3\%$), Brisbane (Eagle R3, 127 qubits, EPLG: $2.2\%$) and Cusco (Eagle R3, 127 qubits, EPLG: $6.8\%$). This 14-qubit experiment tests the potential for a hybrid hardware training. Despite various light error mitigation techniques, the obtained doubly stochastic attention matrices consistently show high entropy (i.e., a tendency toward more uniform distributions), even for larger shot counts. Experimental details and plots are given in Appendix A). As Appendix Figure A1B shows, when comparing to the noise-free ground truth attention matrix, the ordering of the values in the attention matrix was preserved with high precision (spearman $\rho>0.9$ even for moderate shot count). Since the circuit runs within a ViT, successfully preserving the ordering will be key (to not destroy signal). Instead, numerical exactness (cf. Appendix Figure A1A) may be compromised: embedding a noisy quantum attention block (which preserves peak attention scores but also increases entropy) into a Transformer could even be advantageous. The additional entropy may avoid vanishing gradients and act as a form of regularization. This effect is particularly notable compared to its noise-free analog, which remains classically intractable if sufficient qubits are used.